Charlie wants to order a pizza. He can choose to add some or none of the seven toppings from the list below. He can't choose all of them. bacon, mushrooms, olives, pepperoni, pineapple, sausage, tomatoes In how many different ways can Charlie choose the toppings? (The order the toppings are chosen doesn't matter.)
Answer #1

There are total of 7 toppings and for all toppings, we have 2 options: Either to include or not but he can't choose all of them. Hence,

Total possibilities = 27 - 1 = 128 - 1 = 127